# Quantum Spin Liquids in Weak Mott Insulators with a Spin-Orbit Coupling

### v1

@@ Abstract
The weak Mott insulating regime of the triangular lattice Hubbard model exhibits a rich magnetic phase diagram as a result of the ring exchange interaction in the spin Hamiltonian. These phases include the Kalmeyer-Laughlin type chiral spin liquid (CSL) and a valence bond solid (VBS). A natural question arises regarding the robustness of these phases in the presence of a weak spin-orbit coupling (SOC). In this study, we derive the effective spin model for the spin-orbit coupled triangular lattice Hubbard model in the weak Mott insulting regime, including all SOC-mediated spin-bilinears and ring-exchange interactions. We then construct a simplified spin model keeping only the most relevant SOC-mediated spin interactions. Using infinite density matrix renormalization group (iDMRG) we show that the CSL and VBS phases of the triangular lattice Hubbard model can be stabilized in the presence of a weak SOC. The stabilization results from a compensation between the Dzyaloshinskii-Moriya interaction and a SOC-mediated ring exchange interaction. We also provide additional qualitative arguments to intuitively understand the compensation mechanism in the iDMRG quantum phase diagrams. This mechanism for stabilization can potentially be useful for the experimental realization of quantum spin liquids.
@@ I Introduction
Another novel route to achieve frustration via competing interactions is to utilize the charge fluctuations in weak Mott insulators [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. In this case, the virtual charge fluctuations can lead to significant multi-spin interactions because of the small charge gap. For example, starting from the Hubbard model at half-filling, the strong coupling expansion in terms of $t/U$ [35, 36], where $t$ and $U$ are the hopping and on-site repulsion, leads to four-spin ring-exchange interactions that compete with the nearest and further neighbor two-spin interactions. The Hubbard model on the triangular lattice with a moderate charge gap [37, 38, 39, 40, 41, 42, 43] and its effective spin models including ring-exchange interactions [44, 45, 40] have been extensively studied using DMRG. In the $J_{1}$-$J_{\text{ring}}$ model [44], it was found that tuning the ratio of $J_{\text{ring}}/J_{1}$ between 00-$0.4$ changes the ground state successively from a $120^{\circ}$ order, to a Kalmeyer-Laughlin chiral spin liquid (CSL) [46, 47], valence bond solid (VBS), and finally into a zig-zag ordered state. The phase diagram of the $J_{1}$-$J_{2}$-$J_{3}$-$J_{\text{ring}}$ spin model [45] also shows a similar structure when parameterized by $t/U$.
@@ II.1 Hubbard Model with a SOC
$H=\sum_{ij\alpha\beta}\tilde{t}_{ij,\alpha\beta}c^{\dagger}_{i\alpha}c_{j\beta}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (1)
@@ II.2 Effective Spin Hamiltonian in the weak Mott insulating regime
and $S$ is computed order by order in $\tilde{t}/U$ where $S^{(n)}\propto(\tilde{t}/U)^{n}$. At each order, $S^{(n)}$ is chosen such that $H_{\text{eff}}$ does not have any term which takes it out of the singly occupied sector. $S$ is computed up to $\mathcal{O}(\tilde{t}^{3}/U^{3})$, and the strong-coupling expansion in $\tilde{t}/U$ in Eq. (4) is truncated to $\mathcal{O}(\tilde{t}^{4}/U^{3})$. (See Appendix A for details of the canonical transformation). Finally, the resulting Hamiltonian is projected onto the singly-occupied subspace of the triangular lattice to obtain the effective spin model. For simplicity we have also considered $\mathbf{v}_{ij}=v_{z;i,j}\hat{\mathbf{z}}$ (out of plane). This choice is reflected in the symmetry of the resulting effective spin Hamiltonian in Eq. (5), which has a $U(1)$ symmetry. Further, we have considered a uniform nearest neighbour hopping $t$, and a uniform $v_{z}$ with its positive direction on the triangular lattice as defined in Fig. 1(c). Along its positive (negative) direction $\mathbf{v}$ points along $\hat{\mathbf{z}}$ ($-\hat{\mathbf{z}}$). This choice of directions was motivated by studies on organic charge-transfer salts [59, 60, 61, 62] with the space group C2/c that have a DM interaction due to inversion symmetry breaking.
@@
$+\sum_{\langle ij\rangle}D_{z}{S}_{i}^{[x}{S}_{j}^{y]}+H_{\text{ring}}+H_{\text{ring,SOI}},$ (5)
@@ III Quantum Phase Diagrams
Figure 2: (a) Quantum Phase diagram obtained from iDMRG by starting from a CSL phase at $J_{1}=1$, $J_{2}=J_{3}=0.01$, $J_{r}=0.2$ ($t/U=0.097$). The SOC generates the nearest neighbour DM interaction $D_{z}$ and a SOC-mediated ring exchange interaction $J_{r_{1}}$. The two types of $120^{\circ}$ orders in (a) are distinguished by their handedness $\omega$. The $D_{z}$ term prefers the $120^{\circ}_{-}$ phase with $\omega<0$ and the $J_{r_{1}}$ term prefers $120^{\circ}_{+}$ phase with $\omega>0$. The compensation between these opposing tendencies leads to the stability of the CSL phase in a narrow diagonal region, where their combined effect is minimized. Along the vertical axis $J_{r_{1}}/J_{r}=0$, the CSL phase persists until $D_{z}/J_{1}=0.0125$, as in inset. Observables along the vertical dashed line at $J_{r_{1}}/J_{r}=0.4$ are shown in (b)-(f). The CSL has a nonzero scalar chirality $\chi$, a high entanglement entropy $S_{E}$, and a weak intensity at $S(K)$, compared to the $120^{\circ}_{\pm}$ phases.
@@
A CSL is a type of QSL that spontaneously breaks time-reversal and parity symmetries [46, 64]. At the mean-field level the CSL can be described using a fermionic parton construction [65, 1, 5], where spins fractionalize into neutral spin $1/2$ particles. These fermionic partons are in bands with non-zero Chern numbers and are coupled to an emergent $U(1)$ gauge field. The $U(1)$ gauge field is gapped by a Chern-Simons term, and this leads to non-trivial mutual statistics [47, 64] of the deconfined spinon excitations in the CSL. Additionally, the CSL carries propagating chiral edge modes [66], similar to quantum Hall systems [46, 64].
Due to spontaneous time-reversal symmetry breaking in the CSL state, the scalar chirality acquires a non-zero expectation value. The scalar chirality averaged over all triangles is defined as
@@
at the $K$ points (see Figs. 1 (b), 10 & 11) in the Brillouin zone. Classically, the $120^{\circ}$ order has a $3$-site unit cell, and the spins are co-planar having a relative angle of $120^{\circ}$ between each other. Unlike the classical phase, the quantum $120^{\circ}$ phase is highly entangled, but has similar long-ranged correlations as the classical $120^{\circ}$. To classify the two different types of $120^{\circ}$ phases in Fig. 2, we further subdivide the quantum $120^{\circ}$ phase into the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phase based on their handedness. The handedness $\omega$ is defined as:
@@
Figure 3: (a) Quantum Phase diagram obtained from iDMRG by starting from a VBS phase at $J_{1}=1$, $J_{2}=J_{3}=0.012$,$J_{r}=0.24$ ($t/U=0.105$). Similar to Fig. 2, the compensation between the DM interaction $D_{z}$ and SOC mediated-ring exchange $J_{r_{1}}$ leads to the VBS phase being stabilized in the diagonal region. In addition to the $120^{\circ}_{\pm}$ orders, there are narrow regions of the CSL accompanying the VBS. Additionally, a magnetic ordered (MO) phase with long-range correlations, distinct from the $120^{\circ}_{-}$ is identified. Observables along the vertical dashed line $J_{r_{1}}/J_{r}=0.3$ are shown in (b)-(f). The VBS is characterized by a weak intensity at $S(K)$ and strong peaks in the dimer structure factor $D_{3}(M^{\prime\prime})$. To clearly distinguish the CSL around $D_{z}/J_{1}\sim 0.055$ from its neighbouring states, a zoomed-in plot is provided in Fig. 7.
@@
where $D^{n}_{i}=\mathbf{S}_{i}\cdot\mathbf{S}_{i+\mathbf{a}_{n}}$ is the dimer operator along the $\mathbf{a}_{n}$ direction (see Fig. 1 (a)). The peak in the static dimer structure factor occurs at the $M$ point corresponding to the translation vectors of the singlet covering bonds. However, a VBS state obtained from iDMRG is peaked at $M$ points in the dimer structure factor along all three directions $\mathbf{a}_{n}$. This is because the ground state obtained from iDMRG is a superposition of different types of singlet coverings. The dominant singlet covering can be identified from the positions of the peaks and their relative intensities [45].
@@
First, we characterize the nature of the ground state in the absence of a SOC when $t/U=0.105$. From Eq. (8), this value corresponds to $J_{1}=1$, $J_{2}=0.01$, $J_{3}=0.01$, $J_{r}=0.24$. With this parameterization, we find a VBS ground state, in agreement with [45]. This VBS has preferred bonding axes along $\mathbf{a}_{2}$ and $\mathbf{a}_{3}$ in the triangular lattice (see Figs. 1 (a) & 9).
@@
In the presence of a SOC, we find from Fig. 3 (a) that the VBS phase is stabilized along the diagonal region as the strengths of $D_{z}$ and $J_{r_{1}}$ are increased. This stabilization can be attributed to the same mechanism which is a result of compensation between the DM interaction and SOC-mediated ring exchange interactions as discussed in Sec. III.1. Interestingly in Fig. 3 (a), we also observe two bands of the CSL phase, above and below the VBS phase. These CSL bands are stabilized along a diagonal region by a similar mechanism. However, states in these two CSL bands have opposite signed handedness $\omega$. We find that increasing the strength of $J_{r_{1}}$ drives a VBS state into a CSL, we provide a qualitative explanation of this fact in Sec. IV.3. Apart from observing the long-range ordered $120^{\circ}_{-}$ and $120^{\circ}_{+}$ phases in similar regimes compared to Fig. 2 (a), we also observe another magnetically ordered (MO) phase in Fig. 3 (a). Even though the MO phase is long-range ordered, it has a smaller correlation length than the $120^{\circ}_{-}$ (see Figs. 11 & 12). Additionally, the MO phase has a higher entanglement entropy $S_{E}$, and a higher intensity at $D_{3}(M^{\prime\prime})$ compared to the the $120^{\circ}_{-}$ phase. We are unable to fully characterize the MO phase based on the available data from the iDMRG simulations. The MO phase may be related to the AFM-II phase identified in a variational Monte Carlo study [62] of the $J_{1}$-$J_{r}$-$D_{z}$ model. The AFM-II phase in the study [62] is predicted to have peaks in the static spin structure factor close to but not exactly at the $K$ point. However, we could not resolve a similar feature for the MO phase from our iDMRG simulations.
@@ IV.1 Stability of the chiral spin liquid and valence bond solid phases
    and since the spins are coplanar we have used $\left\langle S^{x}_{k}S^{x}_{l}+S^{y}_{k}S^{y}_{l}\right\rangle=\mathbf{S}_{k}\cdot\mathbf{S}_{l}=|S|^{2}\cos(2\pi/3)=-1/8$. The overall negative sign in $H_{\text{eff,B}}$ is consistent with the stability of the $120^{\circ}_{+}$ phase in this regime. This decoupling occurs for 4 different rhombuses having the bond $(i,j)$, leading to the factor of $-1/2$ in $H_{\text{eff,B}}$ (see Fig. 5 (a)).
Now that both limits are written in the form of a DM interaction, it is easy to identify a compensation between $H_{\text{eff,A}}$ and $H_{\text{eff,B}}$ when $H_{\text{eff,A}}+H_{\text{eff,B}}\approx 0$, that is $D_{z}/J_{r_{1}}=0.5$. Comparing with the iDMRG results from Figs. 2 & 3, we find that in the region of stability $D_{z}/J_{r_{1}}\approx 0.44$ (for the CSL) and $D_{z}/J_{r_{1}}\approx 0.52$ (for the VBS) respectively, and are within $12\%$ of that expected from the qualitative argument. Additional evidence in support of this stabilization mechanism can be found by examining the expectation value $\left\langle H_{D_{z}}+H_{J_{r_{1}}}\right\rangle$ at various locations in the phase diagram. In Figs. 4 (a) & (b), we observe the absolute value of this quantity to be minimized and close to zero in the region of stability of the CSL and VBS phases respectively. This observation further demonstrates the two SOC-mediated terms, $H_{D_{z}}$ and $H_{J_{r_{1}}}$, compensate for each other in the region of stability of these phases.

### v2

@@ Abstract
The weak Mott insulating regime of the triangular lattice Hubbard model exhibits a rich magnetic phase diagram as a result of the ring exchange interaction in the spin Hamiltonian. These phases include the Kalmeyer-Laughlin type chiral spin liquid (CSL) and a valence bond solid (VBS). A natural question arises regarding the robustness of these phases in the presence of a weak spin-orbit coupling (SOC). In this study, we derive the effective spin model for the spin-orbit coupled triangular lattice Hubbard model in the weak Mott insulting regime, including all SOC-mediated spin-bilinears and ring-exchange interactions. We then construct a simplified spin model keeping only the most relevant SOC-mediated spin interactions. Using infinite density matrix renormalization group (iDMRG) we show that the CSL and VBS phases of the triangular lattice Hubbard model can be stabilized in the presence of a weak SOC. The stabilization results from a compensation between the Dzyaloshinskii-Moriya interaction and a SOC-mediated ring-exchange interaction. We also provide additional qualitative arguments to intuitively understand the compensation mechanism in the iDMRG quantum phase diagrams. This mechanism for stabilization can potentially be useful for the experimental realization of quantum spin liquids.
@@ I Introduction
Another novel route to achieve frustration via competing interactions is to utilize the charge fluctuations in weak Mott insulators [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. In this case, the virtual charge fluctuations can lead to significant multi-spin interactions because of the small charge gap. For example, starting from the Hubbard model at half-filling, the strong coupling expansion in terms of $t/U$ [35, 36], where $t$ and $U$ are the hopping and on-site repulsion, leads to four-spin ring-exchange interactions that compete with the nearest and further neighbor two-spin interactions. The Hubbard model on the triangular lattice with a moderate charge gap [37, 38, 39, 40, 41, 42, 43, 44, 45] and its effective spin models including ring-exchange interactions [46, 47, 40] have been extensively studied using DMRG. In the $J_{1}$-$J_{\text{ring}}$ model [46], it was found that tuning the ratio of $J_{\text{ring}}/J_{1}$ between 0-$0.4$ changes the ground state successively from a $120^{\circ}$ order, to a Kalmeyer-Laughlin chiral spin liquid (CSL) [48, 49], valence bond solid (VBS), and finally into a zig-zag ordered state. The phase diagram of the $J_{1}$-$J_{2}$-$J_{3}$-$J_{\text{ring}}$ spin model [47] also shows a similar structure when parameterized by $t/U$.
@@ II.1 Hubbard Model with a SOC
$H=-\sum_{ij\alpha\beta}\tilde{t}_{ij,\alpha\beta}c^{\dagger}_{i\alpha}c_{j\beta}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (1)
@@ II.2 Effective Spin Hamiltonian in the weak Mott insulating regime
and $S$ is computed order by order in $\tilde{t}/U$ where $S^{(n)}\propto(\tilde{t}/U)^{n}$. At each order, $S^{(n)}$ is chosen such that $H_{\text{eff}}$ does not have any term which takes it out of the singly occupied sector. $S$ is computed up to $\mathcal{O}(\tilde{t}^{3}/U^{3})$, and the strong-coupling expansion in $\tilde{t}/U$ in Eq. (4) is truncated to $\mathcal{O}(\tilde{t}^{4}/U^{3})$. (See Appendix A for details of the canonical transformation). Finally, the resulting Hamiltonian is projected onto the singly-occupied subspace of the triangular lattice to obtain the effective spin model. For simplicity we have also considered $\mathbf{v}_{ij}=v_{z;i,j}\hat{\mathbf{z}}$ (out of plane). This choice is reflected in the symmetry of the resulting effective spin Hamiltonian in Eq. (5), which has a $U(1)$ symmetry. Further, we have considered a uniform nearest neighbour hopping $t$, and a uniform $v_{z}$ with its positive direction on the triangular lattice as defined in Fig. 1(c). Along its positive (negative) direction $\mathbf{v}$ points along $\hat{\mathbf{z}}$ ($\hat{\mathbf{z}}$). This choice of directions was motivated by studies on organic charge-transfer salts [61, 62, 63, 64] with the space group C2/c that have a DM interaction due to inversion symmetry breaking.
@@
$+\sum_{\langle ij\rangle}D_{z}{S}_{i}^{[x}{S}_{j}^{y]}+H_{\text{ring}}+H_{\text{ring,SOC}},$ (5)
@@ III Quantum Phase Diagrams
Figure 2: (a) Quantum Phase diagram obtained from iDMRG by starting from a CSL phase at $J_{1}=1$, $J_{2}=J_{3}=0.01$, $J_{r}=0.2$ ($t/U=0.097$). The SOC generates the nearest neighbour DM interaction $D_{z}$ and a SOC-mediated ring exchange interaction $J_{r_{1}}$. The two types of $120^{\circ}$ orders in (a) are distinguished by their handedness $\omega$. The $D_{z}$ term prefers the $120^{\circ}_{-}$ phase with $\omega<0$ and the $J_{r_{1}}$ term prefers $120^{\circ}_{+}$ phase with $\omega>0$. The compensation between these opposing tendencies leads to the stability of the CSL phase in a narrow diagonal region, where their combined effect is minimized. Along the vertical axis $J_{r_{1}}/J_{r}=0$, the CSL phase persists until $D_{z}/J_{1}=0.0125$, as in inset. Observables along the vertical dashed line at $J_{r_{1}}/J_{r}=0.4$ are shown in (b)-(f). The CSL has a nonzero scalar chirality $\chi$, a high entanglement entropy $S_{E}$, and a weak intensity at $S(K)$, compared to the $120^{\circ}_{\pm}$ phases. Location of the peaks in the second derivative of the energy in (d), can also be used to identify the location of the phase boundaries.
@@
A CSL is a type of QSL that spontaneously breaks time-reversal and parity symmetries [48, 66]. At the mean-field level the CSL can be described using a fermionic parton construction [67, 1, 5], where the electrons fractionalize into neutral spin $1/2$ particles. These fermionic partons are in bands with non-zero Chern numbers and are coupled to an emergent $U(1)$ gauge field. The $U(1)$ gauge field is gapped by a Chern-Simons term, and this leads to non-trivial mutual statistics [49, 66] of the deconfined spinon excitations in the CSL. Additionally, the CSL carries propagating chiral edge modes [68], similar to quantum Hall systems [48, 66].
Due to spontaneous time-reversal symmetry breaking in the CSL state, the scalar chirality acquires a non-zero expectation value [69]. The scalar chirality averaged over all triangles is defined as
@@
at the $K$ points (see Figs. 1 (b), 11 & 12) in the Brillouin zone. Classically, the $120^{\circ}$ order has a $3$-site unit cell, and the spins are co-planar having a relative angle of $120^{\circ}$ between each other. Unlike the classical phase, the quantum $120^{\circ}$ phase is highly entangled, but has similar long-ranged correlations as the classical $120^{\circ}$. To classify the two different types of $120^{\circ}$ phases in Fig. 2, we further subdivide the quantum $120^{\circ}$ phase into the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phase based on their handedness. The handedness $\omega$ is defined as:
@@
Figure 3: (a) Quantum Phase diagram obtained from iDMRG by starting from a VBS phase at $J_{1}=1$, $J_{2}=J_{3}=0.012$$J_{r}=0.24$ ($t/U=0.105$). Similar to Fig. 2, the compensation between the DM interaction $D_{z}$ and SOC mediated-ring exchange $J_{r_{1}}$ leads to the VBS phase being stabilized in the diagonal region. In addition to the $120^{\circ}_{\pm}$ orders, there are narrow regions of the CSL accompanying the VBS. Additionally, a magnetic ordered (MO) phase with long-range correlations, distinct from the $120^{\circ}_{-}$ is identified. Observables along the vertical dashed line $J_{r_{1}}/J_{r}=0.3$ are shown in (b)-(f). The VBS is characterized by a weak intensity at $S(K)$ and strong peaks in the dimer structure factor $D_{3}(M^{\prime\prime})$. To clearly distinguish the CSL around $D_{z}/J_{1}\sim 0.055$ from its neighbouring states, a zoomed-in plot is provided in Fig. 8.
@@
where $D^{n}_{i}=\mathbf{S}_{i}\cdot\mathbf{S}_{i+\mathbf{a}_{n}}$ is the dimer operator along the $\mathbf{a}_{n}$ direction (see Fig. 1 (a)). The peak in the static dimer structure factor occurs at the $M$ point corresponding to the translation vectors of the singlet covering bonds. However, a VBS state obtained from iDMRG is peaked at $M$ points in the dimer structure factor along all three directions $\mathbf{a}_{n}$. This is because the ground state obtained from iDMRG is a superposition of different types of singlet dimer coverings. The dominant dimer covering can be identified from the positions of the peaks and their relative intensities [47].
@@
First, we characterize the nature of the ground state in the absence of a SOC when $t/U=0.105$. From Eq. (8), this value corresponds to $J_{1}=1$, $J_{2}=0.01$, $J_{3}=0.01$, $J_{r}=0.24$. With this parameterization, we find a VBS ground state, in agreement with [47]. The peaks of the dimer structure factor at $D_{2}(M)$ and $D_{3}(M^{\prime})$ have the highest intensity (see Figs. 1 (a) & 10). This implies the VBS state is, to leading order, a superposition of two types of dimer coverings [47]. These two dimer coverings have their singlet bonding axes oriented along $\mathbf{a}_{2}$ and $\mathbf{a}_{3}$ respectively, in the triangular lattice (see Fig. 4).
Figure 4: The two equal-weight dominant dimer coverings in the VBS phase. The singlet bonding axes in the dimer coverings are along $\mathbf{a}_{2}$ and $\mathbf{a}_{3}$ respectively.
@@
In the presence of a SOC, we find from Fig. 3 (a) that the VBS phase is stabilized along the diagonal region as the strengths of $D_{z}$ and $J_{r_{1}}$ are increased. This stabilization can be attributed to the same mechanism which is a result of compensation between the DM interaction and SOC-mediated ring exchange interactions as discussed in Sec. III.1. Interestingly in Fig. 3 (a), we also observe two bands of the CSL phase, above and below the VBS phase. These CSL bands are stabilized along a diagonal region by a similar mechanism. However, states in these two CSL bands have opposite signed handedness $\omega$. We find that increasing the strength of $J_{r_{1}}$ drives a VBS state into a CSL, we provide a qualitative explanation of this fact in Sec. IV.3. Apart from observing the long-range ordered $120^{\circ}_{-}$ and $120^{\circ}_{+}$ phases in similar regimes compared to Fig. 2 (a), we also observe another magnetically ordered (MO) phase in Fig. 3 (a). Even though the MO phase is long-range ordered, it has a smaller correlation length than the $120^{\circ}_{-}$ (see Figs. 12 & 13). Additionally, the MO phase has a higher entanglement entropy $S_{E}$, and a higher intensity at $D_{3}(M^{\prime\prime})$ compared to the the $120^{\circ}_{-}$ phase. We are unable to fully characterize the MO phase based on the available data from the iDMRG simulations. The MO phase may be related to the AFM-II phase identified in a variational Monte Carlo study [64] of the $J_{1}$-$J_{r}$-$D_{z}$ model. The AFM-II phase in the study [64] is predicted to have peaks in the static spin structure factor close to but not exactly at the $K$ point. However, we could not resolve a similar feature for the MO phase from our iDMRG simulations.
@@ IV.1 Stability of the chiral spin liquid and valence bond solid phases
    and since the spins are coplanar we have used $\left\langle S^{x}_{k}S^{x}_{l}+S^{y}_{k}S^{y}_{l}\right\rangle=\mathbf{S}_{k}\cdot\mathbf{S}_{l}=|S|^{2}\cos(2\pi/3)=-1/8$. The overall negative sign in $H_{\text{eff,B}}$ is consistent with the stability of the $120^{\circ}_{+}$ phase in this regime. This decoupling occurs for 4 different rhombuses having the bond $(i,j)$, leading to the factor of 1/2-1/2 in $H_{\text{eff,B}}$ (see Fig. 6 (a)).
Now that both limits are written in the form of a DM interaction, it is easy to identify a compensation between $H_{\text{eff,A}}$ and $H_{\text{eff,B}}$ when $H_{\text{eff,A}}+H_{\text{eff,B}}\approx 0$, that is $D_{z}/J_{r_{1}}=0.5$. Comparing with the iDMRG results from Figs. 2 & 3, we find that in the region of stability $D_{z}/J_{r_{1}}\approx 0.44$ (for the CSL) and $D_{z}/J_{r_{1}}\approx 0.52$ (for the VBS) respectively, and are within $12\%$ of that expected from the qualitative argument. Additional evidence in support of this stabilization mechanism can be found by examining the expectation value $\left\langle H_{D_{z}}+H_{J_{r_{1}}}\right\rangle$ at various locations in the phase diagram. In Figs. 5 (a) & (b), we observe the absolute value of this quantity to be minimized and close to zero in the region of stability of the CSL and VBS phases, respectively. This observation further demonstrates the two SOC-mediated terms, $H_{D_{z}}$ and $H_{J_{r_{1}}}$, compensate for each other in the region of stability of these phases.